In the scenarios listed below, find the direction
Answer #1

1. Posive charge in electric fielf

F = qE , hereF is in the direction of E

Ans. to the Right

2. -ve charge in electric fielf

F = qE , here F is in opposite direction of E

Ans. to the left

3.Postive charge moving out in magnetic field upward

F = q(vXB) = q (v j x B k)

=qvb i

Ans. Force is to the right

4. -ve charge moving out in magnetic field upward

F = -q(vXB)

Ans. Force is to the left

5. +ve charge moving downward electric field point out of the screen,

F=qE

Direction force is same as electric field.

Ans. Points out of the screen.

6.-ve charge moving downward electric field point out of the screen,

F=-qE

Direction force is opposite as electric field.

Ans. into of the screen.